two resistances 60.36 and 30.09 are connected in parallel. The equivalent resistance is

Grade:12

1 Answers

Arun
4 years ago
Dear Mahesh
 
In parallel combination
 
1/Req = 1/R1 + 1/R2
 
1/Req = 1/60.36 + 1/30.09
 
1/Req = 0.017 + 0.033 = 0.05
 
hence Req = 1/0.05 = 20